# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v2

@@ 4.3 Optimization
4.3.1 Prompt Engineering (PE)
Prompt engineering leverages the inherent capabilities of large language models without additional fine-tuning. By carefully crafting input prompts, practitioners can guide the model to produce desirable embeddings tailored to specific tasks. Several notable methods have been developed. PromptEOL [166] utilizes in-context learning to enhance sentence embedding performance. Echo [185] populates individual texts into a prompt template twice and obtains the textual embedding from the second occurrence position of the text. PromptReps [180] extends the use of prompt engineering to generate retrieval-specific embeddings. By inputting documents and queries along with relevant prompts into the LLM, it utilizes the embedding at the <EOS> position as a feature for retrieval tasks. MetaEOL [171] introduces an unsupervised text embedding technique known as Meta-Task Prompting with Explicit One-Word Limitation. In the study on PromptSTH/SUM [172], the authors introduce two novel prompt engineering techniques—Pretended Chain of Thought and Knowledge Enhancement—which enhance the expressiveness of LLM embeddings and are effective across diverse LLMs. Further, BeLLM [174] explores the impact of backward dependencies in LLMs and then proposes a backward dependency enhanced large language model (BeLLM) that transforms certain attention layers from unidirectional to bidirectional. This modification allows the model to learn more nuanced sentence embeddings, improving performance across various STS tasks and demonstrating the value of incorporating backward dependencies.
@@ 4.3.3 Supervised Contrastive Learning (SCL)
Hard Negatives
Numerous studies on contrastive learning have demonstrated that the quality of negative samples significantly impacts the final training performance of contrastive learning [155, 210, 211]. ANCE [212] explores a two-stage training method. First, it utilizes non-embedding model methods, such as BM25, to extract a batch of negative samples for model training. Subsequently, it employs the trained embedding model to re-mine higher-quality negative samples for further model training. This method has become a standard practice for training embedding models. Conan-Embedding [213] employs a dynamic hard negative mining method to maximize the utilization of high-quality negative examples. However, real-world datasets often suffer from incomplete positive sample labeling due to cost constraints; for most datasets, only around one positive sample is labeled per query. This necessitates careful consideration of the influence of false negatives during negative sample mining. Many researchers suggest that sampling from the middle of the embedding retrieval results or filtering based on similarity scores can effectively enhance model performance [214, 215].
@@
Instruction tuning in embedding models involves adding explicit instructions to the input text to specify downstream tasks and domains or to distinguish whether the input belongs to the query or document side of a retrieval task. This allows the model to differentiate between various inputs and enhance the performance of downstream tasks. Instructor [67] first proposed the method of instruction tuning for embedding models and constructed an instruction tuning dataset containing different tasks such as classification, retrieval, and clustering. TART [51] focused on applying the instruction tuning method to multiple retrieval tasks, creating an instruction tuning training set, BERRI, which covers 40 distinct retrieval tasks. The strong instruction-following capabilities of large language models (LLMs) have further facilitated the exploration of instruction tuning in embedding models based on LLMs.  [153] explored utilizing LLMs like GPT-4 to construct instruction tuning datasets and trained a decoder-only pre-trained model for text embedding. Min et al. [216] introduced a curriculum learning approach for multitask instruction tuning in embeddings. InBedder [191] introduced a novel training methodology. Furthermore, many researchers have focused on building test benchmarks to evaluate different models’ instruction-following abilities and retrieval performance, such as InstructIR [217], followIR [154], and C-STSB [218].
Adaptation Pre-training for Embedding Task
@@
4.3.4 Multi-Task Learning
In the supervised training phase, in addition to the contrastive learning objective, some research explores the introduction of other learning objectives to enhance the final model’s representation capabilities or to maintain the generative abilities of the LLMs, thereby enabling broader applications in downstream tasks. Piccolo2 [221] employs different loss functions for various tasks; for instance, it uses contrastive loss (CL) for retrieval tasks and margin loss for semantic textual similarity (STS) tasks. CDE [222] proposes two methods for creating contextualized document embeddings that improve neural retrieval by incorporating neighboring document information during encoding, resulting in better performance than traditional bi-encoders, particularly in out-of-domain contexts. Furthermore, GritLM [165] simultaneously trains LLMs for both contrastive learning and Next Token Prediction (NTP) tasks, enabling the model to represent text effectively while retaining its original text generation capabilities.
4.3.5 Multi-Stage Training
During the supervised learning phase, numerous studies have explored multi-stage contrastive learning training methods to enhance further the generalization and versatility of the final embedding models. For instance, influenced by earlier work from the BERT era, models like GTR [175], GTE-Qwen2 [192], and BMRetriever [178] have investigated two-stage training strategies that combine weakly supervised contrastive learning (WCL) with supervised contrastive learning (SCL). Weakly supervised contrastive learning primarily leverages a large amount of weakly supervised relevance data collected from public domains (e.g., neighboring text spans and question-answer pairs from the QA community) for training. Although this data may contain noise, the extensive domain coverage provided by typically over a billion data points significantly improves the model’s domain generalization capabilities. Additionally, NV-Retriever proposes a phased supervised training approach tailored to different tasks (e.g., retrieval and STS tasks). In the first stage, NV-Retriever [149] utilizes only retrieval-supervised data with in-batch negatives alongside mined hard negatives. In the second stage, it blends data for retrieval tasks with datasets from other tasks.
@@ 4.4 Commercial Service
Table 5: The Overview of Text Embedding Understanding With LLMS. LM and CD represent Language Model and Context Disillation, separately. PC, ICL, RAG represent Prompt Compression, In-Context Learning and Retrieval-Augmented Generation.
@@
Embedding Source LLM Scale Form Paradigm Projector Dataset
@@
GEIA [237] SRoBERTa, SimCSE, ST5, MPNet GPT-2 133K-209K X CG TR-AI,TR-EI
Vec2Text [238] GTR, OpenAI API T5 5M X PCG 1-Layer MLP TR-EI
M-Vec2Text [239] T5 ME5 3M-5M X PCG 1-Layer MLP TR-EI
@@
Text Revealer [241] BERT, Tiny-BERT GPT-2 - X CG - TR-EI
Tranfer Attack [242] OpenAI API, SBERT, ST5 GPT-base - X PCG 1-Layer MLP TR-EI
5 Text Embedding Understanding with LLMs
LLMs have a powerful paraphrase capability that can be quickly aligned with and interpreted in a variety of already learned embedding spaces of image [243], item [244, 245] and concept [246]. Thus, for different purposes, existing work attempts to make an understanding of text embeddings with the help of LLMs. There are two embedding-related tasks in NLP, long context compression (ICC) and embedding inversion (EI).
@@ 5.1 Methods for Long Context Compression
Long Context Compression proposes that the input for conditioning a language model can be condensed into a smaller, specifically chosen set of words or weights. These methods reduce context length, leading to faster inference speeds and smaller storage requirements. Based on the form of compressed context, this method can be classified into two categories: soft prompt 5.1.1 and context distillation 5.1.2. It is worth noting that in natural language processing, handling long contexts has always been a significant issue. Some works [247, 248, 249, 250, 251, 252, 253] enhance the transformer’s ability to model long contexts by storing and utilizing past representations.  [254] proposed compressing the raw activations of long contexts. These works primarily aim to expand the context of Transformers rather than compressing contexts, but their methods might also be adapted for context compression in the future.
@@ 5.1.1 Soft Prompt
[255] propose soft prompt as a trainable, parameterized prompt to adapt pre-trained language models to specific downstream tasks in natural language processing. In prompt compression, the general idea of this method is to align the soft prompt with the representation of the original long text, thereby enabling the replacement of the original long text with the soft prompt in practical applications. Different methods propose different optimization objectives to obtain sufficiently short soft prompts that can preserve the semantic meaning of the original text as much as possible. According to the training tasks, they can be divided into three categories: alignment, prediction, and restoration.
@@
The prediction method takes the language model task of predicting the next word as the optimization objective.  [229] adapt LLMs into AutoCompressors by compressing long contexts into summary vectors. During training, summary vectors are trained using an unsupervised approach, where long documents are segmented and summary vectors from previous segments are incorporated into language modeling. During inference, all summary vectors are concatenated to form the soft prompt for the long text. [230] extend standard text-to-text transformer models to representation+text-to-text models by transferring contexts to shorter embeddings. [131] adopts a two-stage training strategy to enable a text encoder to compress a long context into a single input token embedding for an LLM. [231] adopts an LLM to transform the token embedding of contexts to shorter embeddings. [232] combines contrastive learning and language model loss to compress and select demonstrations simultaneously.
@@ 5.1.2 Context Distillation
The selective method relies on the capabilities of the LLM itself to directly score each token. [130] employs a small language model to iteratively select the most important tokens by calculating perplexity. [258] also utilizes a small language model and compute self-information for each token. Nugget [259] proposes a text representation method for an encoder-decoder structure that maps a sequence of tokens to a shorter one by utilizing a trained scorer to select top-k tokens in the last layer of the encoder. Nugget2D [259] extends Nugget to the decoder-only LLMs like LLaMA by utilizing selected tokens in all layers of transformers rather than only in the last layer. This method uses a small number of tokens to represent the semantic information of context and generally does not require additional annotated data for training. However, this method cannot produce complete natural language paragraphs that humans can understand, leading to a lack of interpretability. [260] utilizes supervisory signal generated using by LLM to optimize a language model scoring system through reinforcement learning. [261] removes tokens with low cross-attention scores within the context of two documents.
@@ 6.1.1 False Negative detection
Under the embedding training framework of contrastive learning, the model aims to learn appropriate embeddings by pulling positive sample pairs closer while pushing negative sample pairs farther apart. The appearance of false negatives disrupts this learning process. Samples that originally belong to the same category (positive sample relationship) are wrongly regarded as negative samples to participate in the training. This causes the model to receive incorrect "supervisory signals", resulting in the learned embeddings failing to accurately reflect the real semantic relationships among samples. An intuitive idea to solve the false negative problem is to correct the wrongly labeled samples through accurate annotations. However, whether it is the traditional manual annotation method or annotation with the help of large-scale language models (LLMs), both face the challenge of high costs. For manual annotation, in the face of massive data, it is almost an impossible task to expect to annotate thousands of samples for each query. Manual annotation is not only time-consuming and labor-intensive but also easily affected by subjective factors of annotators and issues such as the consistency of annotation standards. On the other hand, although LLMs have demonstrated powerful capabilities in many fields such as natural language processing, using them for annotation also comes with relatively high cost expenditures, including the cost of calling APIs, possible errors in generated results, and limitations on the accuracy of annotating domain-specific knowledge. Therefore, relying on large-scale and accurate annotations to completely eliminate false negatives has significant obstacles in reality.
@@ 7.1 Methods for Cross-Lingual & Cross-Modal Domain
Advanced LLMs already support over a hundred languages [300] and can understand the information in multiple modalities such as vision [301], audio [302], and even physical fields [303]. The powerful understanding of LLMs has been leveraged to represent multi-lingual and multi-modal data.

### v3

@@ 4.3 Optimization
4.3.1 Training Free (TF)
Some works leverage the inherent capabilities of large language models without additional fine-tuning.
Prompt Engineering
By carefully crafting input prompts, practitioners can guide the model to produce desirable embeddings tailored to specific tasks. In the era of PLM, PromptBERT [217] first proposed a prompt-based method to enhance the text representation of the BERT model in training-free setting. Specifically, this method uses a fixed template (e.g., "[X] is [mask]", where [X] is a placeholder.) and leverages the hidden states at the mask position as the sentence representation. Recently, two lines of methods for LLMs have been developed:
  • Summary-Style Prompt. Several approaches leverage prompt engineering to condense input text into a single word, aiming to capture core semantics. For instance, PromptEOL [170] employs the template ‘‘The sentence [X] means in a word:’’ to elicit a one-word summary, where [X] is replaced with the input text. The final hidden state is then used as the sentence embedding. Furthermore, PromptEOL enhances performance by utilizing in-context learning, generating one-word summaries with GPT-4 for a sentence and prepending them to the input as contextual examples. Building on this, PromptReps [184] expands the utility of prompt engineering by not only crafting dense text embeddings but also enabling sparse bag-of-words representations derived from language models. To further enhance LLM embedding expressiveness, PromptSTH/SUM [176] proposes a Pretended Chain of Thought and Knowledge Enhancement. Pretended Chain of Thought simulates step-by-step reasoning by prepending prompts with phrases like "After thinking step by step," encouraging detailed sentence representation. Knowledge Enhancement integrates human summarization expertise by incorporating relevant knowledge into prompts, guiding the model to focus on essential information. MetaEOL [175] guides LLMs to produce embeddings through a series of carefully designed prompts that address multiple representational aspects, then use an average of these meta-task-derived embeddings as the final representation. Similarly, GenEOL [191] introduces a novel approach by generating diverse sentence variations of the target text using an LLM. Subsequently, it computes the final sentence embedding by averaging the embeddings of these generated variants. Different from these, MoEE [192] leverages the inherent structure of Mixture-of-Experts (MoE) LLMs, combining routing weights and hidden states to produce robust embeddings without requiring explicit fine-tuning.
  • Repetition-Style Prompt. The inherent causal attention mechanism in decoder-only Large Language Models (LLMs) can restrict information flow for embedding models, prompting research into mitigation strategies via prompt design. Echo et al. [190] introduce a self-repetition prompt, Rewrite the sentence:[x], rewritten sentence:[x], where [x] represents the input sequence. This method leverages the repetition of the input to enhance contextual understanding, extracting text embeddings through mean pooling applied solely to the second occurrence of [x]. ReBA [193] extends this concept by developing an algorithm that directly modifies the model’s attention matrix, using the repeated input to refine hidden state representations at each position, leading to enhanced performance compared to basic repetition. Token Prepending [177] proposes a plug-and-play, training-free approach. This technique involves prepending the decoded sentence embedding from each layer to the input of the subsequent layer. By doing so, it allows earlier tokens to access the complete sentence embedding, effectively circumventing the limitations imposed by causal attention.
@@ 4.3.3 Supervised Contrastive Learning (SCL)
Prompt-Based Learning
Prompt-based learning [218] aligns the fine-tuning form of downstream tasks with that of the pre-training task, enabling the mode learned during pre-training to be fully utilized. Some prompt-based training-free methods, such as PromptEOL [170], PromptReps [184] and Echo [190], Further, BeLLM [178] explores the impact of backward dependencies in LLMs and then proposes a backward dependency enhanced large language model (BeLLM) that transforms certain attention layers from unidirectional to bidirectional. This modification allows the model to learn more nuanced sentence embeddings, improving performance across various STS tasks and demonstrating the value of incorporating backward dependencies.
@@
General-purpose embedding models confront inherent challenges stemming from the diverse and often conflicting objectives across various downstream tasks. For instance, a sentence pair deemed semantically similar may prove irrelevant in a retrieval context if it fails to address a specific query. Early approaches attempted to mitigate this issue by differentiating symmetric and asymmetric retrieval through the use of prefixes, such as ’query:’ and ’passage:’. However, with the expanding landscape of applications, including classification, clustering, retrieval, and semantic similarity assessment, instruction-based prompting has emerged as a promising strategy. This approach aims to enhance task-specific performance by explicitly incorporating task descriptions within the input.
Instructor [67] firstly pioneered instruction tuning for embedding models, building upon GTR [179] and curating a dataset encompassing various tasks. Building upon this foundation, Data-CUBE [219] advanced the field by implementing curriculum learning within multitask instruction tuning. TART [51] specifically targeted retrieval tasks, constructing the BERRI dataset, which covers 40 distinct retrieval scenarios. The robust instruction-following capabilities of large language models (LLMs) have catalyzed the development of LLM-based embedding models with instruction tuning. InBedder [197] proposes a novel methodology for training instruction-following sentence embedders by exploiting the generative capacity of LLMs. Rather than directly manipulating embedding vectors, InBedder fine-tunes LLMs on abstractive question answering, extracting sentence embeddings from the averaged hidden states of generated answers during inference. E5-Mistral [157] utilizes a large, LLM-generated, instruction-rich training dataset to tune LLM embeddings. Notably, this model initially applied instructions only 0to the query side, a practice subsequently adopted by numerous subsequent works [182, 183, 153, 187, 188, 168, 169, 198, 199, 200, 174, 202, 203]."
In-Context Tuning
The practice of in-context tuning [220], which augments training inputs with contextual examples, has proven effective in enhancing a model’s understanding of contextual dependencies. Following this trend, recent research [187, 202] has explored the adaptation of in-context tuning for LLM-based embedding models. These approaches involve fine-tuning pre-trained LLMs with in-context examples that share semantic similarity with the target input, resulting in improved performance on various downstream tasks."
4.3.4 Multi-Task Learning
Beyond the standard contrastive learning objective, recent research explores the integration of auxiliary learning objectives during supervised training to enhance representation quality and preserve generative capacities in Large Language Models (LLMs). This multi-objective approach aims to broaden the applicability of LLMs across diverse downstream tasks.
Piccolo2 [221] adopts task-specific loss functions, employing contrastive loss (CL) for retrieval and margin loss for semantic textual similarity (STS). This task-adaptive loss strategy optimizes performance for distinct modalities. Similarly, Contextual Document Embeddings (CDE) [222] leverages neighboring document information to generate contextualized embeddings, significantly improving neural retrieval, particularly in out-of-domain scenarios, compared to traditional bi-encoders.
GritLM [169] demonstrates that simultaneous training with contrastive learning and Next Token Prediction (NTP) enables LLMs to achieve robust text representation while maintaining their native generative abilities. This dual-objective approach balances representational and generative demands. ChatRetriever [183] further refines LLMs for retrieval by integrating contrastive learning with masked instruction tuning on high-quality conversational data, thereby enhancing complex session understanding.
DEBATER [188] introduces a continuous thinking process to augment dense retrieval. This method generates a sequence of tokens representing a chain of thought prior to producing the final query representation. Document similarity is then determined by the maximum similarity score between any token in the chain and the document. Training incorporates both contrastive learning and self-distillation to refine the representation.
ULLME [201] proposes Generation-Representation Learning (GRL), a fine-tuning technique that jointly optimizes contrastive learning and generation objectives. GRL minimizes the Kullback-Leibler (KL) divergence between similarity scores derived from learned representations and probability distributions from generation, ensuring internal consistency between the representational and generative aspects of the model.
O1-Embedder [189] utilizes a single LLM for both query expansion and text embedding. The query expansion is modelled as a next token prediction task, where a large-size LLM is used to generate the supervised queries. At inference time, the embedding of the original query and O1-Embedder’s newly generated query are pooled equally as the final query embedding.
4.3.5 Multi-Stage Training
Pretext Task $\rightarrow$ Contrastive Learning
@@ 4.3.5 Multi-Stage Training
Multi-Stage Contrastive Learning
During the supervised learning phase, numerous studies have explored multi-stage contrastive learning training methods to enhance further the generalization and versatility of the final embedding models. For instance, influenced by earlier work from the BERT era, models like GTR [179], GTE-Qwen2 [198], and BMRetriever [182] have investigated two-stage training strategies that combine weakly supervised contrastive learning (WCL) with supervised contrastive learning (SCL). Weakly supervised contrastive learning primarily leverages a large amount of weakly supervised relevance data collected from public domains (e.g., neighboring text spans and question-answer pairs from the QA community) for training. Although this data may contain noise, the extensive domain coverage provided by typically over a billion data points significantly improves the model’s domain generalization capabilities. Additionally, NV-Retriever proposes a phased supervised training approach tailored to different tasks (e.g., retrieval and STS tasks). Specifically, NV-Retriever [153] uses only retrieval-supervised data with in-batch negatives alongside mined hard negatives for the first stage while blending data for retrieval tasks with datasets from other tasks for the second stage.
@@ 4.4 Commercial Service
5 Text Embedding Understanding with LLMs
LLMs have a powerful paraphrase capability that can be quickly aligned with and interpreted in a variety of already learned embedding spaces of image [227], item [228, 229] and concept [230]. Thus, for different purposes, existing work attempts to make an understanding of text embeddings with the help of LLMs. There are two embedding-related tasks in NLP: long context compression (ICC) and embedding inversion (EI).
Table 5: The Overview of Text Embedding Understanding With LLMs. The paradigm part is shortened as follows: Language Model (LM), Context Distillation (CD), Auto Encoding (AE), Context Generation (CG) and Progressive Context Generation (PCG). The task part is shortened as follows: Prompt Compression (PC), In-Context Learning (ICL), Retrieval-Augmented Generation (RAG), Information Retrieval (IR), Attribute Inference (AI), and Embedding Inversion (EI).
@@ 5 Text Embedding Understanding with LLMs
Embedding Source LLM Scale Form Paradigm Projector Task
@@
GEIA [243] SRoBERTa, SimCSE, ST5, MPNet GPT-2 133K-209K X CG AI, EI
Vec2Text [244] GTR, OpenAI API T5 5M X PCG 1-Layer MLP EI
M-Vec2Text [245] T5 ME5 3M-5M X PCG 1-Layer MLP EI
@@
Text Revealer [247] BERT, Tiny-BERT GPT-2 - X CG - EI
Tranfer Attack [248] OpenAI API, SBERT, ST5 GPT-base - X PCG 1-Layer MLP EI
@@ 5.1 Methods for Long Context Compression
Long Context Compression proposes that the input for conditioning a language model can be condensed into a smaller, specifically chosen set of words or dense vectors. These methods reduce context length, leading to faster inference speeds and smaller storage requirements. Based on the form of compressed context, this method can be classified into two categories: soft prompt 5.1.1 and context distillation 5.1.2. Long-context understanding has always been a significant issue in NLP. Some works [249, 250, 251, 252, 253, 254, 255] enhance the transformer’s ability to model long contexts by storing and utilizing past representations. [256] proposed compressing the raw activations of long contexts. These works primarily aim to expand the context of Transformers rather than compressing contexts, but their methods might also be adapted for context compression in the future.
@@ 5.1.1 Soft Prompt
Prompt tuning [257] introduces the trainable, parameterized soft prompt to adapt pre-trained language models to specific downstream tasks in natural language processing. In prompt compression, the general idea of this method is to align the soft prompt with the representation of the original long text, thereby enabling the replacement of the original long text with the soft prompt in practical applications. Different methods propose different optimization objectives to obtain sufficiently short soft prompts that can preserve the semantic meaning of the original text as much as possible. According to the training tasks, they can be divided into three categories: alignment, prediction, and restoration.
Alignment-Based Methods
@@
Prediction-Based Methods
The prediction method takes the language model task of predicting the next word as the optimization objective. [235] adapt LLMs into AutoCompressors by compressing long contexts into summary vectors. During training, summary vectors are trained using an unsupervised approach, where long documents are segmented and summary vectors from previous segments are incorporated into language modeling. All summary vectors are concatenated during inference to form the soft prompt for the long text. [236] extend standard text-to-text transformer models to representation+text-to-text models by transferring contexts to shorter embeddings. [135] adopts a two-stage training strategy to enable a text encoder to compress a long context into a single input token embedding for an LLM. [237] adopts an LLM to transform the token embedding of contexts to shorter embeddings. [238] combines contrastive learning and language model loss to compress and select demonstrations simultaneously.
Restoration-Based Methods
@@ 5.1.2 Context Distillation
Selection-Based Methods
The selective method relies on the capabilities of the LLM itself to directly score each token. [134] employs a small language model to iteratively select the most important tokens by calculating perplexity. [260] also utilizes a small language model and computes self-information for each token. Nugget [261] proposes a text representation method for an encoder-decoder structure that maps a sequence of tokens to a shorter one by utilizing a trained scorer to select top-k tokens in the last layer of the encoder. Nugget2D [261] extends Nugget to the decoder-only LLMs like LLaMA by utilizing selected tokens in all layers of transformers rather than only in the last layer. This method uses a small number of tokens to represent the semantic information of context and generally does not require additional annotated data for training. However, this method cannot produce complete natural language paragraphs that humans can understand, leading to a lack of interpretability. [262] utilizes supervisory signal generated using by LLM to optimize a language model scoring system through reinforcement learning. [263] removes tokens with low cross-attention scores within the context of two documents.
Generation-Based Methods
@@ 6.1.1 False Negative detection
Under the embedding training framework of contrastive learning, the model aims to learn appropriate embeddings by pulling positive sample pairs closer while pushing negative sample pairs farther apart. The appearance of false negatives disrupts this learning process. Samples that originally belonged to the same category (positive sample relationship) are wrongly regarded as negative samples to participate in the training. This causes the model to receive incorrect "supervisory signals", resulting in the learned embeddings failing to accurately reflect the real semantic relationships among samples. An intuitive idea to solve the false negative problem is to correct the wrongly labeled samples through accurate annotations. However, whether it is the traditional manual annotation method or annotation with the help of large-scale language models (LLMs), both face the challenge of high costs. For manual annotation, in the face of massive data, it is almost an impossible task to expect to annotate thousands of samples for each query. Manual annotation is not only time-consuming and labor-intensive but also easily affected by subjective factors of annotators and issues such as the consistency of annotation standards. On the other hand, although LLMs have demonstrated powerful capabilities in many fields such as natural language processing, using them for annotation also comes with relatively high cost expenditures, including the cost of calling APIs, possible errors in generated results, and limitations on the accuracy of annotating domain-specific knowledge. Therefore, relying on large-scale and accurate annotations to completely eliminate false negatives has significant obstacles in reality.
@@ 7.1 Methods for Cross-Lingual & Cross-Modal Domain
Advanced LLMs and MLLMs already support over a hundred languages [302] and can understand the information in multiple modalities such as vision [303], audio [304], and even physical fields [305]. The powerful understanding of LLMs and MLLMs has been leveraged to represent multi-lingual and multi-modal data.